# Mamba-based Efficient Spatio-Frequency Motion Perception for Video Camouflaged Object Detection

### v1

@@ III-D Dual Domain Motion Perception
Frequency-based Long-range Motion Perception. As shown in Fig. 6, FLMP follows a similar process to SLMP. In addition, to further explore the potential of frequency in VCOD, as shown in Fig. 1, we apply FFT to obtain amplitude and phase maps of VCOD images, and then use iFFT to reconstruct amplitude-only and phase-only images in spatial domain. As we can see, amplitude map can represent appearance features such as texture and color in VCOD images, while the phase map encodes image layout and structural information such as object shape and edges. Directly leveraging frequency variations to perceive motion may introduce background distraction which is susceptible in VCOD.
The phase map can mitigate this by filtering out appearance distractions and enhancing structural perception, making it ideal for motion representation. Therefore, we only use phase domain as carrier for motion perception. Specifically, the spatial features $\left\{F_{i}\right\}_{i=1}^{N},F_{i}\in\mathbb{R}^{C\times H\times W}$ are converted to phase maps $\left\{F_{i}^{{}_{Pha}}\right\}_{i=1}^{N},F_{i}^{Pha}\in\mathbb{R}^{C\times H\times W}$ and amplitude maps $\left\{F_{i}^{{}_{Amp}}\right\}_{i=1}^{N},F_{i}^{Amp}\in\mathbb{R}^{C\times H\times W}$ via FFT. The phase maps are then flattened and temporally stacked into spatio-frequency map $M_{f}\in\mathbb{R}^{C\times N\times HW}$. Four-directional sequential modeling is performed on $M_{f}$ to extract phase motion features $\left\{F_{i}^{{}_{Pm}}\right\}_{i=1}^{N},F_{i}^{Pm}\in\mathbb{R}^{C\times H\times W}$. This can be formulated as:
@@
While phase motion features $F_{i}^{Pm}$ provide precise structural cues, amplitude maps $F_{i}^{Amp}$ contain essential image details. Complete frequency representation requires the combination of both. Therefore, we use $F_{i}^{Pm}$ to guide $F_{i}^{Amp}$, filtering out background interference before combining. Specifically, they are input into a Mamba-based fusion module (MFM) with the same structure as SFMF, making the amplitude maps depend on phase-based motion features at both global and local levels. The phase-guided amplitude maps $\left\{F_{i}^{g}\right\}_{i=1}^{N},F_{i}^{g}\in\mathbb{R}^{C\times H\times W}$ are then obtained by applying a sigmoid function. Finally, $F_{i}^{g}$ and $F_{i}^{Pm}$ are combined to yield spatio-frequency motion features $\left\{F_{i}^{fre}\right\}_{i=1}^{N},F_{i}^{fre}\in\mathbb{R}^{C\times H\times W}$ via iFFT. The whole process can be formulated as:
@@ III-E Space and Frequency Motion Fusion
Subsequently, the two sequences are reshaped via cross-merge into a concatenated 2D feature map shaped as $\mathbb{R}^{C\times 2H\times W}$, which is further processed by a linear layer and FFN to encode and exchange spatial features. Finally, a $\times 3$ 3 convolution integrates the concatenated features into the final unified motion representations of shape $\mathbb{R}^{C\times H\times W}$.
@@ III-F Model Optimization
TABLE I: QUANTITATIVE COMPARISON WITH the 14 STATE-OF-THE-ART METHODS FOR VCOD ON TWO BENCHMARKS. BEST RESULTS ARE HIGHLIGHTED IN BOLD. “ $\uparrow$ / $\downarrow$ ” INDICATES THE LARGER OR SMALLER IS BETTER.
Method Year Input MACs MoCA-MASK CAD2016
@@
SINet [35] 2020-CVPR Image 58.18G 0.574 0.185 0.655 0.030 0.221 0.156 0.601 0.204 0.589 0.089 0.289 0.209
SINet-v2 [4] 2021-TPAMI Image 30.70G 0.571 0.175 0.608 0.035 0.211 0.153 0.544 0.181 0.546 0.049 0.170 0.110
ZoomNet [50] 2022-CVPR Image 238.75G 0.582 0.211 0.536 0.033 0.224 0.167 0.587 0.225 0.594 0.063 0.246 0.166
BGNet [51] 2022-IJCAI Image 146.13G 0.590 0.203 0.647 0.023 0.225 0.167 0.607 0.203 0.666 0.089 0.345 0.256
FEDERNet [42] 2023-CVPR Image 90.06G 0.555 0.158 0.542 0.049 0.192 0.132 0.607 0.246 0.725 0.061 0.361 0.257
FSPNet [52] 2023-CVPR Image 708.28G 0.594 0.182 0.608 0.044 0.238 0.167 0.539 0.220 0.553 0.145 0.309 0.212
PUENet [53] 2023-TIP Image 407.90G 0.594 0.204 0.619 0.037 0.302 0.212 0.673 0.427 0.803 0.034 0.499 0.389
RCRNet [54] 2019-ICCV Video 203.14G 0.597 0.174 0.583 0.025 0.194 0.137 - - - - - -
PNS-Net [55] 2021-MICCAI Video 22.18G 0.576 0.134 0.562 0.038 0.189 0.133 0.678 0.369 0.720 0.043 0.409 0.308
MG [56] 2021-ICCV Video 353.1G 0.547 0.165 0.537 0.095 0.197 0.141 0.484 0.314 0.558 0.370 0.351 0.260
SLT-Net [19] 2022-CVPR Video 58.18G 0.656 0.357 0.785 0.021 0.387 0.310 0.679 0.420 0.805 0.033 0.445 0.342
TSP-SAM [20] 2024-CVPR Video 1414.12G 0.673 0.400 0.766 0.012 0.421 0.345 0.681 0.500 0.853 0.031 0.496 0.393
IMEX [21] 2024-TMM Video - 0.661 0.371 0.778 0.020 0.409 0.319 0.684 0.452 0.813 0.033 0.469 0.370
EMIP [22] 2025-TIP Video 79.29G 0.675 0.381 - 0.015 0.426 0.333 0.719 0.514 - 0.028 0.536 0.425
Ours - Video 10.88G 0.684 0.382 0.804 0.010 0.459 0.369 0.729 0.573 0.842 0.034 0.634 0.509
@@ IV-A Datasets and Experiment Settings
Datasets. Following existing VCOD methods [19, 21, 20, 22], the backbone is first pretrained on the COD image dataset COD10K [4]. Subsequently, the entire model is trained on the MoCA-Mask [19] training set, and tested on both the MoCA-Mask test set and the CAD [57] dataset. COD10K is widely used in COD studies [13, 14, 15], containing 3,040 training images which is currently the largest publicly available dataset for COD. The MoCA-MASK dataset includes 87 video sequences, with 71 sequences (19,313 frames) for training and 16 sequences (3,626 frames) for testing. Every fifth frame is manually annotated with bounding box and pixel-level segmentation mask. The camouflaged animal dataset (CAD) is a small dataset of camouflaged animals, comprising 9 video sequences and a total of 836 frames. Pixel-level segmentation masks are manually annotated on every fifth frame. In our experiments, the entire CAD dataset is used as the test set.
[…]
Implementation Details. We implement our model using PyTorch on a NVIDIA GeForce RTX 4090 (24G) GPU. All images are resized to $\times 386$ 386. For both training and pretraining data, we apply data augmentation strategies including random cropping, random flipping, and color enhancement. The model adopts VMamba [32] as the backbone, which is first pretrained on the training set of COD10K. Pretraining takes 8 hours for 100 epochs with a batch size of 6, and backbone parameters are optimized using Adam with a learning rate of 1e-5, reduced by a factor of 10 every 50 epochs. The model is then trained on the MoCA-MASK training set, initializing with the pretrained weights. Training takes 2.5 hours for 10 epochs, using a batch size of 2, and the Adam optimizer with a learning rate of 1e-6.
@@ IV-B Performance Comparisons
TABLE III: ABLATION STUDY OF FREQUENCY LEARNING METHODS IN AFE
@@
IV-C Ablation Studies
[…]
Effectiveness of AFE. Our Mamba-based AFE module innovatively achieves adaptive frequency component enhancement through a frequency-domain sequential scanning strategy. To further validate the effectiveness of this design, we replace the Mamba block in AFE with conventional frequency learning methods such as convolution $\times 3$ 3 and self-attention, and we also replace our sequential scanning strategy with the cross scan of VMamba. The results shown in Table III demonstrate the significant advantages of our method in frequency learning. Furthermore, the proposed sequential scanning strategy outperforms mainstream cross scan in frequency map serialization, as it effectively addresses the issue of semantic incoherence caused by large intensity disparities between frequency components during self-attention’s global modeling, making it more suitable for frequency learning.
TABLE IV: ABLATION STUDY OF LONG-RANGE MOTION PERCEPTION METHODS IN SLMP AND FLMP.
@@ IV-C Ablation Studies
Effectiveness of SLMP and FLMP. This work presents an innovative Mamba-based dual-domain motion perception approach that simultaneously captures spatial and frequency-domain motion cues. As shown in Table IV, we replace SLMP and FLMP with traditional long-range motion modeling methods such as 3D CNN, ConvLSTM [63], and ConvGRU [64]. Experimental results demonstrate that our Mamba-based modules outperform them. Furthermore, as shown in Fig. 9, we visualize the intermediate motion features generated by different long-sequence motion modeling methods through heatmaps. The visualization demonstrates that the motion features modeled by SLMP exhibit more precise target localization and lower background noise interference. Both quantitative and qualitative experimental results clearly demonstrates the advantages of our SSM-based motion perception through spatio-temporal and frequency-temporal scanning strategies.
To reduce interference from spatial appearances such as color and texture, we adopt phase maps as motion carriers instead of scanning frequency map directly. To validate this, we replace phase-guided modeling with direct modeling on frequency map (denoted as B1 in Table IV). Experimental results show that our method performs better, verifying the superiority of phase-guided frequency motion perception.
TABLE V: ABLATION STUDY OF MOTION FEATURE FUSION METHODS IN SFMF.
@@
To further validate the superiority of SFMF, as shown in Table V, we replace SFMF with convolution $\times 3$ 3 and cross-attention, the results show that SFMF achieves better performance than traditional methods. Furthermore, Fig. 11 demonstrates that our SFMF module effectively captures local motion details, and globally integrating cross-domain context to filter irrelevant noise through its point-to-point and sequence-to-sequence fusion strategy.
@@ V CONCLUSION
In this work, we introduce frequency features into VCOD to enhance spatial discriminability and global perception. Furthermore, considering the temporal variation of frequency energy can effectively encode motion, and Mamba enables efficient sequence modeling, we propose a dual-domain motion perception network, Vcamba. Specifically, to alleviate Mamba’s limitation in modeling spatial locality, we introduce RFVSS, which incorporates multi-scale receptive fields. As selective-scan can dynamically focus on key frequency components, we propose AFE. It incorporates a novel sequential scanning strategy to ensure semantic coherence. Then we construct a dual-branch structure comprising SLMP and FLMP to model spatio-temporal and frequency-temporal dependencies for dual-domain motion perception. Finally SFMF integrates spatial and frequency motions globally and locally via our dual-domain sequence concatenation strategy. Extensive experiments show that Vcamba outperforms state-of-the-art methods on two datasets with lower computational cost, confirming its effectiveness and efficiency.

### v2

@@ III-D Dual Domain Motion Perception
Frequency-based Long-range Motion Perception. As shown in Fig. 7, FLMP follows a similar process to SLMP. In addition, to further explore the potential of frequency in VCOD, as shown in Fig. 1, we apply FFT to obtain amplitude and phase spectrums of VCOD images, and then use iFFT to reconstruct amplitude-only and phase-only images in spatial domain. As we can see, amplitude represents appearances such as texture and color, while the phase encodes image layout and structural information. Directly leveraging frequency variations to perceive motion may introduce background distraction which is susceptible in VCOD.
The phase can mitigate this by filtering out appearance distractions and enhancing structural perception, making it ideal for motion representation. In addition, as shown in section III-A , based on the Fourier shift theorem [48], object motion corresponds to phase variation in the Fourier spectrum while leaving amplitude intact. Compared to amplitude, phase exhibits stronger robustness and sensitivity in perceiving such variations.
Therefore, we only use phase spectrum as carrier for motion perception. Specifically, the spatial features $\left\{F_{i}\right\}_{i=1}^{N},F_{i}\in\mathbb{R}^{C\times H\times W}$ are converted to phase spectrum $\left\{F_{i}^{{}_{Pha}}\right\}_{i=1}^{N},F_{i}^{Pha}\in\mathbb{R}^{C\times H\times W}$ and amplitude spectrum $\left\{F_{i}^{{}_{Amp}}\right\}_{i=1}^{N},F_{i}^{Amp}\in\mathbb{R}^{C\times H\times W}$ via FFT. The phase are then flattened and temporally stacked into spatio-frequency map $M_{f}\in\mathbb{R}^{C\times N\times HW}$. Four-directional sequential modeling is performed on $M_{f}$ to extract phase motion features $\left\{F_{i}^{{}_{Pm}}\right\}_{i=1}^{N},F_{i}^{Pm}\in\mathbb{R}^{C\times H\times W}$. This can be formulated as:
@@
While phase motion features $F_{i}^{Pm}$ provide precise structural cues, amplitude $F_{i}^{Amp}$ contain essential image details. A complete frequency representation thus requires combining both. To this end, we employ $F_{i}^{Pm}$ to guide $F_{i}^{Amp}$, filtering out background interference before integration. Specifically, they are fed into a Mamba-based fusion module (MFM) with the same structure as SFMF, allowing amplitude spectrum filtering to be modulated by phase-based motion features at both global and local levels. The phase-guided amplitude maps $\left\{F_{i}^{g}\right\}_{i=1}^{N},F_{i}^{g}\in\mathbb{R}^{C\times H\times W}$ are generated by first applying a sigmoid function to the filtered amplitude features, and then multiplying them with the original amplitude spectrum. Finally, $F_{i}^{g}$ and $F_{i}^{Pm}$ are merged via iFFT to produce spatio-frequency motion features $\left\{F_{i}^{fre}\right\}_{i=1}^{N},F_{i}^{fre}\in\mathbb{R}^{C\times H\times W}$. The entire process can be expressed as:
@@ III-E Space and Frequency Motion Fusion
Subsequently, the two sequences are reshaped via cross-merge into a concatenated 2D feature map shaped as $\mathbb{R}^{C\times 2H\times W}$, which is further processed by a linear layer and FFN to encode and exchange spatial features. Finally, a $3\times 3$ convolution integrates the concatenated features into the final unified motion representations of shape $\mathbb{R}^{C\times H\times W}$.
@@ III-F Model Optimization
TABLE I: Quantitative comparison with the 14 state-of-the-art methods for VCOD on two benchmarks. Best video results are highlighted in BOLD. $\uparrow$ / $\downarrow$” indicates the larger or smaller is better.
Method Year Input Backbone MACs MoCA-MASK CAD2016
@@
SINet [17] 2020-CVPR Image ResNet-50 19.42G 0.574 0.185 0.655 0.030 0.221 0.156 0.601 0.204 0.589 0.089 0.289 0.209
SINet-v2 [15] 2021-TPAMI Image Res2Net-50 12.28G 0.571 0.175 0.608 0.035 0.211 0.153 0.544 0.181 0.546 0.049 0.170 0.110
ZoomNet [49] 2022-CVPR Image ResNet-50 95.50G 0.582 0.211 0.536 0.033 0.224 0.167 0.587 0.225 0.594 0.063 0.246 0.166
BGNet [59] 2022-IJCAI Image Res2Net-50 58.45G 0.590 0.203 0.647 0.023 0.225 0.167 0.607 0.203 0.666 0.089 0.345 0.256
FEDERNet [28] 2023-CVPR Image ResNet-50 18.01G 0.555 0.158 0.542 0.049 0.192 0.132 0.607 0.246 0.725 0.061 0.361 0.257
FSPNet [33] 2023-CVPR Image ViT 141.66G 0.594 0.182 0.608 0.044 0.238 0.167 0.539 0.220 0.553 0.145 0.309 0.212
PUENet [75] 2023-TIP Image ViT 163.16G 0.594 0.204 0.619 0.037 0.302 0.212 0.673 0.427 0.803 0.034 0.499 0.389
RCRNet [67] 2019-ICCV Video ResNet-50 203.14G 0.597 0.174 0.583 0.025 0.194 0.137 - - - - - -
PNS-Net [36] 2021-MICCAI Video Res2Net-50 22.18G 0.576 0.134 0.562 0.038 0.189 0.133 0.678 0.369 0.720 0.043 0.409 0.308
MG [68] 2021-ICCV Video VGG-style 353.1G 0.547 0.165 0.537 0.095 0.197 0.141 0.484 0.314 0.558 0.370 0.351 0.260
SLT-Net [7] 2022-CVPR Video PVT 58.18G 0.656 0.357 0.785 0.021 0.387 0.310 0.679 0.420 0.805 0.033 0.445 0.342
TSP-SAM [35] 2024-CVPR Video SAM 1414.12G 0.673 0.400 0.766 0.012 0.421 0.345 0.681 0.500 0.853 0.031 0.496 0.393
IMEX [34] 2024-TMM Video ResNet-50 - 0.661 0.371 0.778 0.020 0.409 0.319 0.684 0.452 0.813 0.033 0.469 0.370
EMIP [74] 2025-TIP Video PVT 79.29G 0.675 0.381 - 0.015 0.426 0.333 0.719 0.514 - 0.028 0.536 0.425
Vcamba (ours) - Video VMamba 16.49G 0.684 0.382 0.804 0.010 0.459 0.369 0.729 0.573 0.842 0.034 0.634 0.509
@@ IV-A Datasets and Experiment Settings
Datasets. Following existing VCOD methods [7, 34, 35, 74], the backbone is first pretrained on the COD image dataset COD10K [15]. Subsequently, the entire model is trained on the MoCA-Mask [7] training set, and tested on both the MoCA-Mask test set and the CAD [3] dataset. COD10K is widely used in COD studies [27, 6, 42], containing 3,040 training images which is currently the largest publicly available dataset for COD. The MoCA-MASK dataset includes 87 video sequences, with 71 sequences (19,313 frames) for training and 16 sequences (3,626 frames) for testing. Every fifth frame is manually annotated with bounding box and pixel-level segmentation mask. The camouflaged animal dataset (CAD) is a small dataset of camouflaged animals, comprising 9 video sequences and a total of 836 frames. Pixel-level segmentation masks are manually annotated on every fifth frame. In our experiments, the entire CAD dataset is used as a test set.
Metric. Following existing VCOD methods [7, 34, 35, 74], we evaluate Vcamba on six standard evaluation metrics: S-measure($S_{\alpha}$) [14], weighted F-measure($F_{\beta}^{\omega}$) [46], mean Enhanced-alignment measure($E_{\phi}$) [16], mean absolute error($M$) [51], meanDice ($mDice$), meanIoU ($mIoU$). The better performance corresponds to higher $S_{\alpha}$, $F_{\beta}^{\omega}$, $E_{\phi}$, $mDice$, mIoUmIoU, and lower $M$.
Implementation Details. We implement our model using PyTorch on a NVIDIA GeForce RTX 4090 (24G) GPU. All images are resized to $386\times 386$. For both training and pretraining data, we apply data augmentation strategies including random cropping, random flipping, and color enhancement. The model adopts VMamba [44] as the backbone, which is first pretrained on the training set of COD10K. Pretraining takes 8 hours for 100 epochs with a batch size of 6, and backbone parameters are optimized using Adam with a learning rate of 1e-5, reduced by a factor of 10 every 50 epochs. The model is then trained on the MoCA-MASK training set, initializing with the pretrained weights. Training takes 2.5 hours for 10 epochs, using a batch size of 2, and the Adam optimizer with a learning rate of 1e-6.
@@ IV-B Performance Comparisons
IV-C Ablation Studies
Specifically, to assess the complementary role of frequency in improving discriminability of spatial appearance, we conduct a comparative experiment by completely removing the frequency branch, retaining only the spatial branch (denoted as A7 in Table II). In this setting, the model solely relies on spatial-domain features to extract camouflaged cues and perceive motion information. The results reveal a sharp performance decline, with key metrics decreasing by 1.3%($S_{\alpha}$), 4.6%($F_{\beta}^{\omega}$) and 5.1%($E_{\phi}$). This empirically demonstrates that frequency features not only provide complementary discriminative cues but also offer global context and motion information that help preserve object structural integrity. These results also demonstrate that combining spatial and frequency domains enables more comprehensive feature representations.
TABLE III: Ablation study of different backbones.
Method MACsMACs $S_{\alpha}\uparrow$ $F_{\beta}^{\omega}\uparrow$ $mIoU\uparrow$
PVTv2-B5 91.24G 0.676 0.362 0.364
ViT-B 139.59G 0.641 0.304 0.294
Res2Net-50 45.75G 0.584 0.199 0.200
VMamba (Ours) 16.49G 0.684 0.382 0.369
Effectiveness of VMamba. VMamba enables global sequence modeling with linear time complexity, while its selective scanning allows the model to adaptively focus on critical regions. Compared with other visual Mambas, VMamba explicitly adopts a cross-scan mechanism to model long-range dependencies across both horizontal and vertical directions, making it better suited for 2D visual representations. To validate the effectiveness of it, as shown in Table III, we replace the backbone with PVTv2-B5 [62], Res2Net [22] and ViT-B [13], and conduct experiments under identical training strategies. VMamba achieves the best performance across all metrics with lower computational cost, indicating its superior performance–efficiency trade-offs in VCOD.
TABLE IV: ABLATION STUDY OF FREQUENCY LEARNING METHODS IN AFE
@@ IV-C Ablation Studies
Effectiveness of AFE. Our Mamba-based AFE module innovatively achieves adaptive frequency component enhancement through a frequency-domain sequential scanning (FSS) strategy. To further validate the effectiveness of this design, we replace the Mamba block in AFE with conventional frequency learning methods such as convolution $3\times 3$ and self-attention, and we also replace our FSS with the cross scan of VMamba, as it is the most representative and performant scanning strategy among existing Mamba variants. The results shown in Table IV demonstrate the advantages of our method in frequency learning. Furthermore, the proposed FSS outperforms mainstream cross scan in frequency map serialization, as it effectively preserves the continuity, causality, and progressive structural evolution inherent in frequency component superposition after unfolding them into ordered sequences.
TABLE V: ABLATION STUDY OF LONG-RANGE MOTION PERCEPTION METHODS IN SLMP AND FLMP.
@@
Effectiveness of SLMP and FLMP. This work presents an innovative Mamba-based dual-domain motion perception approach that simultaneously captures spatial and frequency-domain motion cues. As shown in Table V, we replace SLMP and FLMP with traditional long-range motion modeling methods such as 3D CNN, ConvLSTM [56], and ConvGRU [2]. Experimental results demonstrate that our Mamba-based modules outperform them. Furthermore, as shown in Fig. 10, we visualize the intermediate motion features generated by different long-sequence motion modeling methods through heatmaps. The visualization demonstrates that the motion features modeled by SLMP exhibit a more precise target localization and less background noise interference. Both quantitative and qualitative experimental results clearly demonstrate the advantages of our SSM-based motion perception through spatio-temporal and frequency-temporal scanning strategies.
TABLE VI: Ablation study of different motion modeling strategies in FLMP.
Method Carrier Guidance $S_{\alpha}\uparrow$ $F_{\beta}^{\omega}\uparrow$ $mIoU\uparrow$
Amplitude-only Amplitude Spectrum 0.679 0.377 0.362
Amplitude + Phase Frequency Spectrum 0.679 0.375 0.358
No Guidance Phase Spectrum 0.680 0.376 0.357
Phase-only (Ours) Phase Spectrum 0.684 0.382 0.369
Effectiveness of Phase-Based Motion Modeling. In the FLMP, we explicitly separate the spectral phase and magnitude. Motion is perceived by modeling temporal variations of phase, which is further used to guide amplitude filtering. As shown in Table VI, we conduct ablation studies comparing phase-only motion modeling with amplitude-only and phase–amplitude combination (i.e., frequency spectrum) strategies. The results indicate that the phase-only strategy achieves the best performance, whereas the other variants suffer performance degradation due to the introduction of additional noise. This observation aligns with the Fourier shift theorem, where phase variations directly correspond to spatial displacement and are more robust to appearance changes. Moreover, directly combining original amplitude with phase leads to performance drops, confirming the necessity of phase-guided amplitude filtering for noise suppression.
TABLE VII: ABLATION STUDY OF MOTION FEATURE FUSION METHODS IN SFMF.
@@
To further validate the superiority of SFMF, as shown in Table VII, we replace SFMF with convolution $3\times 3$ and cross-attention, the results show that SFMF achieves better performance than traditional methods. Furthermore, Fig. 12 demonstrates that our SFMF module effectively captures local motion details, and globally integrating cross-domain context to filter irrelevant noise through its point-to-point and sequence-to-sequence fusion strategy.
IV-D Limitation
While Mamba theoretically reduces computational complexity via linear-time sequence modeling, this advantage does not totally translate into superior runtime performance in practical deployments. Compared to highly optimized self-attention on GPUs, Mamba’s selective scan relies on hardware- and implementation-specific optimizations [25, 12, 1, 44]. Existing general visual Mamba, such as VMamba [44] and Vision Mamba [79], do not exhibit a clear advantage in inference speed (FPS) over well-optimized Transformers. In our experiments, Vcamba achieves 14 FPS with VMamba, comparable to 13 FPS with PVT, suggesting that GPU optimization significantly constrains the practical efficiency of visual Mamba. Further hardware-aware optimizations are therefore necessary to fully exploit its theoretical advantages.
@@ V CONCLUSION
In this work, we propose a dual-domain motion perception network called Vcamba for VCOD that leverages both spatial and frequency cues. Frequency features are introduced to enhance spatial discriminability and global perception. We find that frequency components exhibit a structural evolution pattern during ordered superposition, forming a causal sequence that can be effectively modeled by SSMs. Based on this insight, we introduce an adaptive frequency enhancement (AFE) module with a novel frequency-domain sequential scanning (FSS) strategy to enable effective frequency learning. To further exploit motion cues, we design space-based and frequency-based long-range motion perception (SLMP, FLMP) modules to model spatio-temporal and frequency-temporal dependencies for dual-domain motion perception. A spatial and frequency motion fusion (SFMF) module is finally employed to unify complementary motion cues via our dual-domain sequence concatenation strategy. Extensive experiments show that Vcamba outperforms state-of-the-art methods on two datasets with lower computational cost, confirming its effectiveness and efficiency.